# Microseismic event classification with a lightweight Fourier Neural Operator model

### v1

@@ Abstract
Real-time monitoring of induced seismicity is crucial for mitigating operational hazards, relying on the rapid and accurate classification of microseismic events from continuous data streams. However, while many deep learning models excel at this task, their high computational requirements often limit their practical application in real-time monitoring systems. To address this limitation, a lightweight model based on the Fourier Neural Operator (FNO) is proposed for microseismic event classification, leveraging its inherent resolution-invariance and computational efficiency for waveform processing. In the STanford EArthquake Dataset (STEAD), a global and large-scale database of seismic waveforms, the FNO-based model demonstrates high effectiveness for trigger classification, with an F1 score of 95% even in the scenario of data sparsity in training. The new FNO model greatly decreases the computer power needed relative to current deep learning models without sacrificing the classification success rate measured by the F1 score. A test on a real microseismic dataset shows a classification success rate with an F1 score of 98%, outperforming many traditional deep-learning techniques. A combination of high success rate and low computational power indicates that the FNO model can serve as a methodology of choice for real-time monitoring of microseismicity for induced seismicity. The method saves computational resources and facilitates both post-processing and real-time seismic processing suitable for the implementation of traffic light systems to prevent undesired induced seismicity.
keywords:
@@ Introduction
It is well-documented that the injection or extraction of fluids in subsurface geological formations can trigger induced seismicity [healy1968denver]. Fluid injection typically induces geomechanical instability by reducing the effective normal stress due to increased pore pressure [zoback2010reservoir]. In addition, induced seismicity may also result from other mechanisms, such as fault reactivation triggered by cooling [kivi2022cooling], reservoir compaction [segall1989earthquakes, vlek2019groningen], or stress redistribution [kettlety2020stress]. The diverse mechanisms capable of initiating seismic events, combined with uncertainties regarding subsurface conditions, including stress fields, thermal gradients, and fault geometry, result in limited a priori predictability of induced seismicity regarding its spatial occurrence, timing, and magnitude.
@@
Such hazards are reduced by seismic monitoring with real-time detection and characterization of induced seismicity during anthropogenic operations [verdon2021green]. The seismic event characterization requires triggering, classification of triggers, and phase picking, which is a challenge in both earthquake seismology and microseismic monitoring. Traditionally, seismic analysts have determined earthquake signals by visually inspecting three components of waveforms of multiple stations and manually selecting P- and S-wave phase arrivals. However, this method is time-consuming and prone to errors. In particular, these tasks are difficult for microseismic monitoring due to the typically low magnitudes of microseismic events and their susceptibility to noise interference. Visual detection is impractical in real-time operations, and the growing volume of seismic data makes manual analysis impossible, requiring automated algorithms. The most commonly used algorithms for event detection (i.e., triggering and classification) are STA/LTA [allen1978automatic] and template matching [gibbons2006detection]. Although STA/LTA provides computational efficiency, its sensitivity to low-magnitude events and time-varying noise is limited, and its performance is somewhat subjective as it depends on the choice of STA/LTA parameters. The subjectivity of STA/LTA detection is overcome by selecting low thresholds in STA/LTA detection, resulting in a high rate of triggers with false detections. Hence, the triggers need to be classified into true and false detections. Template matching, on the other hand, is robust in detecting small events, but struggles with waveform variability and high computational requirements, making it difficult to apply in real-time [yoon2015earthquake]. Furthermore, the template matching requires previously recorded events - the templates - rarely available at induced seismicity monitoring.
@@
With the growing interest in the potential of machine learning (ML) in the seismological community, significant efforts have been devoted to improving the capabilities of event identification. The application of ML to seismic event classification and detection has been studied since the 1990s, and many methodologies have been proposed. Many of these approaches use artificial neural networks applied to features extracted from recorded seismic signals [gentili2006automatic, maity2014novel, qu2020automatic]. We note that these algorithms are applied to stationary networks with previously recorded seismicity, which is used for training. With the development and widespread availability of computational resources, deep learning (DL) has become a transformative tool for solving long-standing problems in science and engineering disciplines, including event detection and classification in seismology. In recent years, several deep learning models for detection and phase picking have been published [ross2018generalized, mousavi2019cred, woollam2019convolutional, zhu2019phasenet, soto2021deepphasepick]. Existing deep learning models applied to seismic data analysis differ in several key dimensions, including network architecture, such as convolutional, recurrent, and attention-based frameworks; data input representation, either in the time or frequency domain; and output labeling methods, ranging from discrete point predictions to continuous sequence labeling. In addition, these models vary substantially in complexity and depth, ranging from relatively shallow configurations to significantly deeper architectures.
@@
Various supervised DL methods have been developed, which can be classified based on the type of information they provide, from determining whether an event occurs in a triggered interval of data (classification) to determining the arrival time of the event (picking). Several binary window classification approaches have been developed, in which DL models analyze a time window to predict the presence or absence of an event. For example, Wilkins et al. (2020) developed a custom convolutional neural network (CNN) to detect microseismic events associated with mining operations, achieving a ten-fold increase in event detection compared to human experts [wilkins2020identifying]. Although earlier methods often focused on individual geophones, Shaheen et al. (2021) presented a CNN-based approach to detect small-magnitude events using data from an entire shallow borehole network in Groningen, the Netherlands [shaheen2021groningennet]. In addition, Ross et al. (2018) proposed a generalized phase detection (GPD) method by training a CNN on manually labeled data from the Southern California seismic network to classify data intervals as containing P-waves, S-waves, or noise [ross2018generalized].
@@
Both non-neural and neural network (NN) approaches have shown significant advantages over traditional trace-based detection methods. Chen (2020) presented an unsupervised fuzzy clustering method that utilizes features such as STA/LTA triggers, interval powers and means to achieve robust microseismic event detection even in high-noise environments [chen2020automatic]. Meanwhile, Zheng et al. (2018) and Birnie and Hansteen (2022) used recurrent long short-term memory (LSTM) neural network architectures [zheng2018automatic, birnie2022bidirectional]. Zheng et al. (2018) trained their LSTM model on laboratory-simulated microseismic events and successfully applied it to field data [zheng2018automatic]. Birnie and Hansteen (2022) further improved the LSTM model by incorporating bidirectionality and training it on various synthetic datasets with band-pass noise [birnie2022bidirectional]. Mousavi et al. (2019) developed a CNN-RNN earthquake detector (CRED) that combines convolutional layers with bidirectional LSTM blocks in the residual structure to improve microearthquake detection [mousavi2019cred]. Using a different approach, Birnie et al. (2021) treated detection as a semantic segmentation problem for the array records, where seismic data are represented as images, and the network predicts the presence of an event at each pixel [birnie2021introduction]. More recently, Awais et al. (2023) demonstrated that combining super-virtual refraction interferometry with the unsupervised DBSCAN clustering algorithm enables accurate and automatic first-break traveltime picking even in noisy seismic data sets [awais2023first].
@@
Neural operators, as extensions of neural networks, are very effective in constructing surrogate models for various scenarios governed by partial differential equations (PDEs) [kovachki2023neural]. The Fourier neural operator (FNO) [li2020fourier], in particular, excels at training and generalizing PDEs from data. Unlike conventional numerical solvers that solve a single instance of the PDE, the proposed FNO-based approach provides a resolution-invariant solution to the event detection and classification problem, improving both accuracy and efficiency.
@@
Most deep learning models require a high number of trainable parameters on the order of hundreds of thousands. This, in turn, leads to increasing computational resources and time needed to train such a model. Deploying models with millions of parameters significantly reduces their practicality, particularly in applications demanding real-time data processing, such as microseismic event detection and phase picking. One of our goals in this study is to explore lightweight neural operator models with less computational burden. In the end, reducing the complexity of the computational models leads to less environmental impact, including carbon emissions and pollution.
@@
In this study, we propose a lightweight seismic signal detection model based on FNO. We first apply the FNO framework to classify seismic events from the STanford EArthquake Dataset (STEAD) [mousavi2019stanford]. We train and evaluate FNO on real earthquake datasets, demonstrating high accuracy. Then we train and test our algorithm on real microseismic data with further improved performance. We also benchmark this method with leading DL algorithms widely used in the seismological community [zhu2019phasenet, mousavi2020earthquake]. Showing similar performance with PhaseNet, our model has much fewer trainable parameters, which on a large scale, leads to less computation burden both in training and testing large amounts of seismic data. These results highlight the benefits of using FNO for seismic event detection, paving the way for near real-time and cost-effective microseismic monitoring capabilities.
@@
For this study, we first used data from the STanford EArthquake Dataset (STEAD) [mousavi2019stanford] to analyze seismic events. The dataset is a comprehensive, globally acquired collection of seismic signals covering both local earthquakes and non-earthquake events (seismic noise) classified by humans. It includes approximately 1.2 million three-component seismograms, totaling more than 19,000 hours of seismic recordings. The data were acquired by a network of 2,613 seismic instruments around the world, located within a 350 km radius of recorded earthquakes. Each sample consists of 6,000 data points recorded over one minute at a sampling rate of 100 Hz, and three channels: east-west (E), north-south (N), and vertical (Z). The dataset is carefully labeled, facilitating the development and evaluation of machine learning models for earthquake detection and classification. STEAD serves as a valuable resource for improving automated seismic event analysis and real-time monitoring capabilities for large earthquakes.
This study also utilizes microseismic datasets, which comprise triggered microseismic waveforms recorded during hydraulic fracturing operations. The monitoring array consisted of nine three-component surface seismic sensors distributed across an approximate 100 km² surface area, operating at a sampling frequency of 250 Hz. We downsample the dataset to a 100 Hz sampling rate, as this is the original sampling rate used for all models evaluated here, and the majority of the observed seismic signal is below 50 Hz. Although the triggered dataset contains no false seismic events, some waveform records from individual stations are noisy and lack manual phase picks. The final processed dataset consists of approximately 5,400 waveform segments from approximately 600 manually labeled seismic events, encompassing both local and near seismic events. The near seismic events are typically located at distances greater than 10 km from the nearest recording station.
@@
We evaluate five models for the microseismic event classification task using their original model weights (Table 1). All models are originally trained and tested here on single-station three-component data. The model architectures and weights are available in the SeisBench [woollam2022seisbench].
@@
Orig. train size (# events) 200 440,000 65,700 1,100,000 3,375,000 623,000
@@
The CNN-RNN Earthquake Detector (CRED) is designed exclusively for earthquake detection (i.e., trigger and classification) and does not support seismic phase identification or phase picking [mousavi2019cred]. Inputs into the network are spectrograms of three-component seismograms. CRED analyzes spectrogram representations of seismic waveforms with durations of 30 seconds, sampled at 100 Hz. Its internal architecture integrates convolutional neural network (CNN) layers followed by long short-term memory (LSTM) units, ultimately producing a prediction curve indicating earthquake detections at specific time points within the waveform. For training, earthquake detection labels were originally defined from the onset of the manually picked P-wave arrival, extending to a duration of 2.4 times the interval between P- and S-wave arrivals. In cases where datasets lacked S-wave picks or the detections represented teleseismic P-wave arrivals, detection labels were adjusted to begin at the P-wave arrival and extend for 20 seconds after P-wave arrival. The initial training of CRED involved a dataset comprising 550,000 event waveforms and an equal number of waveforms without signal (only seismic noise) collected from Northern California. In this study, CRED was evaluated solely for the earthquake detection task, utilizing the output of the pick detection function.
@@
DeepPhasePick (DPP) is a suite of models designed for earthquake detection and seismic phase picking [soto2021deepphasepick]. Its detection module employs a convolutional neural network (CNN) structure utilizing depth-wise separable convolutions, which assign probabilities of noise (N), P-phase (P), or S-phase (S) to 5-second waveform time intervals. The intervals to be classified are extracted close to the earthquake origin times and analyst-picked P- and S-phase arrivals, respectively. The model receives these three-component seismic waveform windows as inputs and outputs three-probability vectors corresponding to the predicted class labels (P, S, or N). Originally, DPP was trained on a dataset comprising 25,647 P-phase waveforms, 14,397 S-phase waveforms, and 25,647 noise waveforms collected around the 1995 Mw = 8.1 Antofagasta and the 2007 Mw = 7.7 Tocopilla earthquakes in Northern Chile. Here, we use the DPP trained on the data from the Italian Seismic Network (INSTANCE) [michelini2021instance], as models trained on it perform best on average in all tasks [munchmeyer2022picker].
@@
Earthquake Transformer (EQTransformer) is a neural network designed for joint earthquake detection, seismic-phase identification, and phase (onsets) picking [mousavi2020earthquake]. EQTransformer operates on waveform segments of 60 s duration sampled at 100 Hz, outputting three probability prediction traces corresponding respectively to earthquake detection, P-phase, and S-phase occurrences at each time point within the window. Internally, EQTransformer integrates convolutional neural networks (CNNs), long short-term memory (LSTM) units, and self-attention layers. During training, EQTransformer extensively leverages data augmentation techniques: adding Gaussian noise, inserting random data gaps, and randomly dropping channels. Additionally, EQTransformer performs cyclical time shifts of waveform segments to allow phase picks to appear at arbitrary positions within the window. The labeling of earthquake detection followed the same approach used for CRED. EQTransformer was originally trained on the STEAD dataset [mousavi2019stanford].
@@
Generalized Phase Detection (GPD) is a seismic-phase identification model that analyzes short waveform windows of 4 s duration, sampled at 100 Hz, and classifies each window as P-phase, S-phase, or noise for a single station [ross2018generalized]. Although the model is trained using a 2 Hz high-pass filter, it effectively performs across a wide frequency range, demonstrating that the exact training filter is not strictly required when applying the method to new data. GPD primarily performs phase classification, but applying the trained model in a sliding-window approach enables its use for phase onset detection. In the original setup, seismic arrivals were restricted to occur between seconds 1 and 3 of each input window. For inference, the trained GPD model was applied using a sliding-window approach with no overlap. GPD was originally trained and validated using 4.5 million waveform samples from Southern California, evenly distributed among P-phase arrivals, S-phase arrivals, and noise segments. According to Ross et al. (2018), the probability threshold to distinguish events is set to solve specific problems, for example, to create an earthquake catalog minimizing false positives or not to have any false negatives in earthquake early warning [ross2018generalized].
@@
PhaseNet is a U-Net-based deep neural network specifically designed for seismic phase arrival-time picking [zhu2019phasenet]. It processes waveform segments of 30 seconds sampled at 100 Hz, generating continuous probability traces for P- and S-wave arrivals, each of the same duration as the input window. The threshold of probability is set to 0.5 for both P and S picks. Structurally, PhaseNet employs large convolutional filter sizes, incorporates convolutional layers without stride, and utilizes residual connections. The original PhaseNet model was trained and validated on a dataset comprising 779,514 seismic waveforms with labeled P- and S-wave arrivals from Northern California.
@@
We present a neural operator-based algorithm for the seismic events classification. Fourier Neural Operator (FNO) layers are used to process temporal information using fast Fourier transforms, allowing solutions to be computed quickly. FNO layers are particularly well suited for applications where the domain is a regular grid, which corresponds to the typical structure of seismogram data. Since seismograms are typically sampled at regular time intervals, FNO layers offer an efficient means of processing and encoding this data. A key advantage of neural operators over traditional neural networks is their sample invariance, which allows the input features to be discretized on different grids during each estimation. This eliminates the need for model retraining and allows the algorithm to function efficiently in a variety of seismic recording settings. This property is critical for event detection in dynamic seismic networks, especially for microseismic monitoring, where the acquisition setup may be different for each case study.
We modify the model architecture proposed by Kashefi and Mukerji (2024) for the event classification [kashefi2024novel]. The FNO architecture used in this study consists of three 1-D spectral convolution blocks (Figure 1). The FNO blocks capture both global and local waveform characteristics, allowing for efficient detection of seismic events in complex time series data exhibiting varying signal-to-noise ratios. Our algorithm takes as input a triggered seismic wavefield recorded with three components at a seismic station and outputs a binary result: 1 - signal, 0 - noise.
Figure 1: Schematic graph of the proposed FNO-based framework for seismic event detection. The input is a three-component seismic wavefield, and the output is a binary function: 1 - signal, 0 - noise. (adapted after Kashefi and Mukerji (2024) [kashefi2024novel])
@@
For earthquake detection, we configure the network with 32 modes and a width of 64 in the FNO blocks. The training process uses the AdamW optimizer with a learning rate of $1\times 10^{-4}$ along with the ReduceLROnPlateau scheduler to reduce the learning rate when the validation loss plateaus. The loss function is the binary cross-entropy (BCE), which the model seeks to minimize during training. The seismic events are randomly divided into training (100 one-minute waveforms: 50 signal and 50 noise), validation (same as training), and test (2000 waveforms: 1000 signal and 1000 noise) sets. The validation set is used to fine-tune the learning rate, while the test set is used to evaluate the model’s performance.
We used surface monitoring microseismic dataset for training and validation, and adjusted the neural network architecture to have 10 modes and a width of 20 within the Fourier Neural Operator (FNO) blocks. During model training, the AdamW optimizer was employed with an initial learning rate of $1\times 10^{-3}$, complemented by the same ReduceLROnPlateau scheduler. The dataset was randomly divided into three subsets: a training set of 200 one-minute waveforms (100 seismic signals and 100 noise segments), an identical validation set for tuning hyperparameters, and a larger test set of 1000 waveforms (500 seismic signals and 500 noise segments) to rigorously evaluate the predictive performance of the model.

### v2

# Seismic event classification with a lightweight Fourier Neural Operator model

@@ Abstract
Abstract
Real-time monitoring of induced seismicity is critical to mitigate operational risks, relying on the rapid and accurate classification of triggered data from continuous data streams. Deep learning models are effective for this purpose but require substantial computational resources, making real-time processing difficult. To address this limitation, a lightweight model based on the Fourier Neural Operator (FNO) is proposed for the classification of microseismic events, leveraging its inherent resolution-invariance and computational efficiency for waveform processing. In the STanford EArthquake Dataset (STEAD), a global and large-scale database of seismic waveforms, the FNO-based model demonstrates high effectiveness for trigger classification, with an F1 score of 95% even in the scenario of data sparsity in training. The new FNO model greatly decreases the computer power needed relative to current deep learning models without sacrificing the classification success rate measured by the F1 score. A test on a real microseismic dataset shows a classification success rate with an F1 score of 98%, outperforming many traditional deep-learning techniques. The reduced computational cost makes the proposed FNO model well suited for deployment in resource-constrained, near–real-time seismic monitoring workflows, including traffic-light implementations. The source code for the proposed FNO classifier will be available at: https://github.com/ayratabd/FNOclass.
KEYWORDS
@@ INTRODUCTION
It is well-documented that the injection or extraction of fluids in subsurface geological formations can trigger induced seismicity (healy1968denver). Fluid injection typically induces geomechanical instability by reducing the effective normal stress due to increased pore pressure (zoback2010reservoir). In addition, induced seismicity may also result from other mechanisms, such as fault reactivation triggered by cooling (kivi2022cooling), reservoir compaction (segall1989earthquakes; vlek2019groningen), or stress redistribution (kettlety2020stress). The diverse mechanisms capable of initiating seismic events, combined with uncertainties regarding subsurface conditions, including stress fields, thermal gradients, and fault geometry, result in limited a priori predictability of induced seismicity regarding its spatial occurrence, timing, and magnitude.
@@
Such hazards are reduced by seismic monitoring with real-time detection and characterization of induced seismicity during anthropogenic operations (verdon2021green). The seismic event characterization requires triggering, classification of triggers, and phase picking, which is a challenge in both earthquake seismology and microseismic monitoring. Traditionally, seismic analysts have determined earthquake signals by visually inspecting three components of waveforms of multiple stations and manually selecting P- and S-wave phase arrivals. However, this method is time-consuming and prone to errors. In particular, these tasks are difficult for microseismic monitoring due to the typically low magnitudes of microseismic events and their susceptibility to noise interference. Visual detection is impractical in real-time operations, and the growing volume of seismic data makes manual analysis impossible, requiring automated algorithms. The most commonly used algorithms for event detection (i.e., triggering and classification) are STA/LTA (allen1978automatic) and template matching (gibbons2006detection). Although STA/LTA provides computational efficiency, its sensitivity to low-magnitude events and time-varying noise is limited, and its performance is somewhat subjective as it depends on the choice of STA/LTA parameters. The subjectivity of STA/LTA detection is overcome by selecting low thresholds in the detection, resulting in a high rate of triggers with false detections. Hence, the triggers need to be classified into true and false detections. Template matching, on the other hand, is robust in detecting small events, but struggles with waveform variability and high computational requirements, making it difficult to apply in real-time (yoon2015earthquake). Furthermore, the template matching requires previously recorded events as the templates, which are rarely available at induced seismicity monitoring.
@@
With the growing interest in machine learning (ML) in the seismological community, significant efforts have been devoted to better utilize it in trigger classification. ML has been used in the classification and detection of seismic triggers since the 1990s and many approaches have been suggested. Many of them use artificial neural networks applied to features extracted from recorded seismic signals (gentili2006automatic; maity2014novel; qu2020automatic). We note that these algorithms are applied to stationary networks with previously recorded seismicity, which is used as training data. This is usually not possible for induced seismicity where the networks are newly installed to monitor anthropogenic activity. In recent years, several DL models for detection and phase picking have been published (ross2018generalized; mousavi2019cred; woollam2019convolutional; zhu2019phasenet; soto2021deepphasepick). Existing DL models applied to seismic data analysis differ in several key dimensions, including network architecture, such as convolutional, recurrent, and attention-based frameworks; data input representation, either in the time or frequency domain; and output labeling methods, ranging from discrete point to continuous predictions. These models differ in complexity levels, ranging from relatively simple to deeper architectures.
@@
Various supervised DL methods have been developed, which can be classified based on the type of information they provide, from determining whether an event occurs in a triggered interval of data (classification) to determining the arrival time of the event (picking). Several binary window classification approaches have been developed, in which DL models analyze a time window to predict the presence or absence of an event. For example, wilkins2020identifying developed a custom convolutional neural network (CNN) to detect microseismic events associated with mining operations, achieving a ten-fold increase in event detection compared to human experts. Although earlier methods often focused on individual seismic stations, shaheen2021groningennet presented a CNN-based approach to detect small-magnitude events using data from an entire network of stations (shallow borehole seismic stations in Groningen, the Netherlands). In addition, ross2018generalized proposed a generalized phase detection (GPD) method by training a CNN on manually labeled data from the Southern California seismic network to classify data intervals as containing P-waves, S-waves, or noise. For arrival-time estimation in microseismic settings, transfer-learning and U-Net–type architectures have been used to improve phase picking when labeled microseismic picks are limited (zhang2022phase).
@@
Both non-neural and neural network (NN) approaches have shown significant advantages over traditional trace-based detection methods. chen2020automatic proposed an unsupervised fuzzy clustering method that employs various features such as STA/LTA triggers, interval powers and means to effectively identify microseismic events even when the waveforms are highly noisy. Meanwhile, zheng2018automatic and birnie2022bidirectional used recurrent long short-term memory (LSTM) network architecture. zheng2018automatic trained their LSTM model on laboratory-simulated microseismic events and successfully applied it to field data. birnie2022bidirectional further improved the LSTM model by incorporating bidirectionality and training it on various synthetic datasets with band-pass noise. mousavi2019cred developed a CNN-RNN earthquake detector (CRED) that combines convolutional layers with bidirectional LSTM blocks in the residual structure to improve microearthquake detection. Using a different approach, birnie2021introduction treated detection as a semantic segmentation problem for the array records, where seismic data are represented as images, and the network predicts the presence of an event at each pixel. More recently, awais2023first demonstrated that combining super-virtual refraction interferometry with the unsupervised DBSCAN clustering algorithm enables accurate and automatic first-break traveltime picking even in noisy seismic data sets. Beyond detection and picking, DL methods have increasingly been used for rapid microseismic monitoring tasks such as source localization (often trained on synthetic waveforms and applied to field data) and processing of large distributed acoustic sensing (DAS) datasets (wamriew2021deep; vinard2022localizing; yu2024daseventnet).
@@
Neural operators, as extensions of neural networks, are very effective in constructing surrogate models for various scenarios governed by partial differential equations (PDEs) (kovachki2023neural). The Fourier neural operator (FNO) (li2020fourier), in particular, excels at training and generalizing PDEs from data. Unlike conventional numerical solvers that solve a single instance of the PDE, the proposed FNO-based approach provides a resolution-invariant solution to the event detection and classification problem, improving both accuracy and efficiency.
@@
Most deep learning models require a high number of trainable parameters on the order of hundreds of thousands. This, in turn, leads to increasing computational resources and time needed to train such a model. Models with millions of parameters are impractical, especially for real-time applications such as microseismic event detection and phase picking. One of our goals in this study is to explore lightweight neural operator models with less computational burden. In the end, reducing the complexity of the computational models leads to less environmental impact, including carbon emissions and pollution.
@@
In this study, we propose a lightweight seismic signal detection model based on FNO. We first apply the FNO framework to classify seismic events from the STanford EArthquake Dataset (STEAD) (mousavi2019stanford). We train and evaluate FNO on real earthquake datasets, demonstrating high accuracy. Then we train and test our algorithm on real microseismic data with further improved performance. We also benchmark this method with leading DL algorithms widely used in the seismological community (zhu2019phasenet; mousavi2020earthquake). Showing similar performance with PhaseNet, our model has much fewer trainable parameters, which on a large scale, leads to less computation burden both in training and testing large amounts of seismic data. The results show that the use of FNO for detecting seismic events can be useful for near-real time and cost-effective microseismic monitoring.
@@
For this study, we first used data from the STanford EArthquake Dataset (STEAD) (mousavi2019stanford) to analyze seismic events. STEAD is a large dataset of seismic signals worldwide and encompasses local earthquakes and non-earthquake seismic events (seismic noise) labeled by humans. It includes approximately 1.2 million three-component seismograms, totaling more than 19,000 hours of seismic recordings. The data were acquired by a network of 2,613 seismic instruments around the world, located within a 350 km radius of recorded earthquakes. Each sample consists of 6,000 data points recorded over one minute at a sampling rate of 100 Hz, and three channels: east-west (E), north-south (N), and vertical (Z). The dataset is labeled, which is important for developing and testing machine learning algorithms for identifying and classifying earthquakes.
This study also utilizes a microseismic dataset comprising waveforms recorded by a surface array during hydraulic fracturing operations. The acquisition geometry consisted of nine three-component geophones distributed over an area of approximately 100 km², operating at a sampling frequency of 250 Hz. We downsampled the continuous recordings to 100 Hz, as the dominant frequency of the observed seismic signal is below 50 Hz and 100 Hz is the original sampling rate for all models tested. The dataset contains low-magnitude induced seismicity ($0.5<M<2.5$) with distinct P- and S-wave arrivals, as well as background noise, which includes non-seismic events specific to the field. From the continuous data, we assembled a balanced set of approximately 1,400 60-sec waveforms by selecting approximately equal numbers of event and noise windows (to prevent class imbalance during training). This compilation includes up to 600 manually labeled local and near-network events (epicentral distances > 10 km from the center of network), supplemented by representative noise intervals. All waveforms were preprocessed with amplitude normalization to standardize inputs for the deep-learning models.
We intentionally refrained from applying synthetic data augmentation techniques, such as superimposing multiple events or inserting zero-value gaps, to preserve the authentic statistical properties of the field recordings. While the dataset naturally contains rare instances of overlapping events and near-zero intervals, we prioritized learning from these intrinsic field characteristics rather than artificially synthesizing complex scenarios.
@@
We evaluate five models for the microseismic event classification task using their original model weights (Table 1). All models are originally trained and tested here on single-station three-component data. The model architectures and weights are available in the SeisBench (woollam2022seisbench).
@@
Orig. train size (# events) 200 440,000 32,400 1,100,000 3,375,000 623,000
@@
The CNN-RNN Earthquake Detector (CRED) is designed exclusively for earthquake detection (i.e., trigger and classification) and does not support seismic phase identification or phase picking (mousavi2019cred). Inputs into the network are spectrograms of three-component seismograms. CRED analyzes spectrogram representations of seismic waveforms with durations of 30 seconds, sampled at 100 Hz. Its internal architecture integrates convolutional neural network (CNN) layers followed by long short-term memory (LSTM) units, ultimately producing a prediction curve indicating earthquake detections at specific time points within the waveform. The initial training of CRED involved a dataset comprising 550,000 event waveforms and an equal number of waveforms without signal (only seismic noise) collected from Northern California. In this study, CRED was evaluated solely for the earthquake detection task, utilizing the output of the pick detection function.
@@
DeepPhasePick (DPP) is a suite of models designed for earthquake detection and seismic phase picking (soto2021deepphasepick). Its detection module employs a convolutional neural network (CNN) structure utilizing depth-wise separable convolutions, which assign probabilities of noise (N), P-phase (P), or S-phase (S) to 5-second waveform time intervals. The intervals to be classified are extracted close to the earthquake origin times and analyst-picked P- and S-phase arrivals, respectively. The model receives these three-component seismic waveform windows as inputs and outputs three-probability vectors corresponding to the predicted class labels (P, S, or N). Originally, DPP was trained on a dataset comprising more than 65,000 waveforms collected around the 1995 Mw = 8.1 Antofagasta and the 2007 Mw = 7.7 Tocopilla earthquakes in Northern Chile. Here, we use the DPP trained on the data from the Italian Seismic Network (INSTANCE) (michelini2021instance), as models trained on it perform best on average in all tasks (munchmeyer2022picker). To adapt this architecture to our 60 s recordings, we partitioned each waveform into 12 non-overlapping 5 s intervals and aggregated the predictions, such that the entire window was classified as an event if the probability in any sub-interval exceeded the threshold.
@@
Earthquake Transformer (EQTransformer) is a neural network designed for joint earthquake detection, seismic-phase identification, and phase (onsets) picking (mousavi2020earthquake). The model operates on waveform segments of 60 s duration sampled at 100 Hz, outputting three probability prediction traces corresponding respectively to earthquake detection, P-phase, and S-phase occurrences at each time point within the window. Internally, EQTransformer integrates convolutional neural networks (CNNs), long short-term memory (LSTM) units, and self-attention layers. There are also various data augmentation techniques used for the training data: adding Gaussian noise, random data gaps, and random channel drops. It also shifts waveform segments in cycles, allowing the phase picks to occur at any point in the window. The model was originally trained on the STEAD dataset (mousavi2019stanford). For the binary classification task, we utilized the model’s dedicated detection output channel, classifying a window as a positive event if the detection probability exceeded the evaluation threshold at any sample point within the 60 s trace.
@@
Generalized Phase Detection (GPD) is a seismic-phase identification model that analyzes short waveform windows of 4 s duration, sampled at 100 Hz, and classifies each window as P-phase, S-phase, or noise for a single station (ross2018generalized). GPD primarily performs phase classification, but applying the trained model in a sliding-window approach enables its use for phase onset detection (similarly to DPP). GPD was originally trained and validated using 4.5 million waveform samples from Southern California, evenly distributed among P-phase arrivals, S-phase arrivals, and noise segments.
@@
PhaseNet is a U-Net-based deep neural network specifically designed for seismic phase arrival-time picking (zhu2019phasenet). It processes waveform segments of 30 seconds sampled at 100 Hz, generating continuous probability traces for P- and S-wave arrivals, each of the same duration as the input window. The threshold of probability is set to 0.5 for both P and S picks. Structurally, PhaseNet employs large convolutional filter sizes, incorporates convolutional layers without stride, and utilizes residual connections. The original PhaseNet model was trained and validated on a dataset comprising 779,514 seismic waveforms with labeled P- and S-wave arrivals from Northern California. Similarly to the EQTransformer, we utilized the model’s noise output channel, classifying a window as a positive event if the noise probability fell below the evaluation threshold at any point within the trace.
@@
We propose a neural operator approach for classifying seismic events. This algorithm uses Fourier Neural Operator layers to process temporal data using fast Fourier transforms, allowing solutions to be computed quickly. The FNO blocks encode the seismogram data efficiently due to their regular sampling nature. A key advantage of neural operators over traditional neural networks is their sample invariance, which allows the input features to be discretized on different grids during each estimation. This eliminates the need for model retraining and allows the algorithm to function efficiently in a variety of seismic recording settings. This property is critical for event detection in dynamic seismic networks, especially for microseismic monitoring, where the acquisition setup may be different for each case study.
We modify the model architecture proposed by kashefi2024novel for the seismic event classification. The FNO architecture used in this study consists of three 1-D spectral convolution blocks (Figure 1). The FNO blocks capture both global and local waveform characteristics, allowing for efficient detection of seismic events in complex time series data exhibiting varying signal-to-noise ratios. The input is a three-component seismic waveform, and the output is a single scalar, which is the probability that the window contains a seismic signal.
Figure 1: A simple diagram illustrates the proposed Fourier Neural Operator (FNO) method for detecting seismic events. The input is a three-component seismic waveform window (E, N, Z). The input is first transformed into a higher-dimensional feature space, then processed by three connected 1-D FNO blocks. Each block uses a spectral convolution and a local 1x1 convolution, then a nonlinear activation function, to capture both global and local waveform characteristics. The features are then combined by a global max pooling, and passed through a fully connected classifier that outputs a single number, which is the probability that the window contains a seismic signal (1 = signal, 0 = noise).
@@
The training process uses the AdamW optimizer with a learning rate of $1\times 10^{-3}$ along with the ReduceLROnPlateau scheduler to reduce the learning rate when the validation loss plateaus. The loss function is the binary cross-entropy (BCE), which the model seeks to minimize during training. For earthquake detection, the seismic events are randomly divided into training (100 to 500 one-minute waveforms: balanced signal and noise), validation (50 signal and 50 noise samples), and test (500 signal and 500 noise) sets. The microseismic dataset was randomly divided into three subsets: a training set of 200 one-minute waveforms (100 seismic signals and 100 noise segments), an identical validation set, and a larger test set of 1000 waveforms (500 seismic signals and 500 noise segments). The validation set is used to fine-tune the learning rate, while the test set is used to evaluate the model’s performance.